# How Do Multilingual Models Remember? Investigating Multilingual Factual Recall Mechanisms

### v1

@@ 6 Conclusion
In this paper, we analyzed the mechanisms involved in the recall of factual knowledge in 10 languages using multilingual transformer-based LMs. We compare these findings to previous research on English recall in English autoregressive LLMs. We discovered that some mechanisms, such as the flow of subject representations in the later layers and the extraction phase, are shared between multilingual and monolingual models. However, we also identified notable differences, including the absence of knowledge localization in the early MLPs and the joint role of MLPs and attention modules during the extraction phase in multilingual models. Finally, we also studied when language encoding occurs during the recall process by patching representations. In XGLM, we observed two distinct phases: initially, the relation flows to the last token, followed by the language. In contrast, in mT5, both the relation and language flow to the last token at similar layers. This suggests that while the relation and subject representations are multilingual and can be used to extract objects across languages, the extraction phase itself appears to be language-specific, as the language is encoded prior to extraction. These findings provide new evidence to understand the factual knowledge recall in transformer LMs, contributing with analyzes in different architectures and languages, and suggesting new directions for the study of cross-lingual transfer.
@@ 7 Limitations
In this paper, we examined two model architectures, leaving out the effects of model sizes, instruction fine-tuning, or models like Llama that can behave multilingually but have less coverage and less multilingual pre-training data. Additionally, our analyses were conducted on 500-1000 examples per language, which we believe provides a sufficient sample size for generalization; however, the results are inherently limited by the relations present in the mParaRel dataset, which may not capture all factual nuances. Additionally, although we analyzed 10 diverse languages, many more languages exist, and further research is needed to confirm the generalizability of our findings across a broader linguistic spectrum. Lastly, we described the main mechanisms found in XGLM and mT5, however other weaker mechanisms could be at play, which could describe, for example, the low extraction rates found for some languages (Figure 16) or the few examples where the object seems to be encoded from early layers before the extraction takes place (Figure 31).
@@ Appendix A Related Work
Factual Knowledge Recall.
Petroni et al. (2019) analyzed the factual knowledge encoded in pre-trained language models, leading to further multilingual research. [^17] One may object to calling the ability to solve cloze tests as knowledge. LLMs are sometimes inconsistent and cannot always justify the propositions they generate. We side with Fierro et al. (2024) in thinking knowledge may nevertheless be the appropriate concept, since LLMs are mostly consistent and can sometimes provide justification, e.g., in virtue of world models or training data attribution. Kassner et al. (2021); Jiang et al. (2020) extended this work by translating the LAMA dataset (Petroni et al., 2019) and evaluating the factual knowledge acquired by multilingual pre-trained models, while Wang et al. (2024) evaluated it from the perspective of knowledge edits (Zhu et al., 2021). Building on this, Yin et al. (2022) curated a dataset that focuses on the cultural aspect of factual knowledge. While Fierro and Søgaard (2022) presented the mParaRel dataset181818mParaRel is a machine-translated version of the English templates from ParaRel (Elazar et al., 2021), where subjects and objects are translated using WikiData identifiers. to examine the consistency of factual predictions in multilingual pre-trained language models. In a related effort, Qi et al. (2023) measured the cross-lingual consistency of multilingual LLMs by evaluating the consistency across languages, rather than just across paraphrases in the same language.
Interpretability on Factual Knowledge Recall.
@@
Activation Patching.
Ghandeharioun et al. (2024) introduced the Patchscope framework to study what an intermediate representation encodes. In their experiments, they focused on “decoding” the contents of a latent representation by patching it into the forward pass of another input. Similarly, Zhang and Nanda (2024) explored the differences and implementation nuances of patching representations versus adding Gaussian noise. More relevant to the type of patching we conducted in this work, Dumas et al. (2024) investigated a word translation task to assess whether the model encodes concepts that are disentangled from the language during translation. This builds on earlier findings by Wendler et al. (2024), who observed that models like Llama-2 (Touvron et al., 2023) seem to use English as an internal pivot language. Consistent with our results, Dumas et al. (2024); Foroutan et al. (2022) provided evidence for the existence of language-agnostic representations.
@@
In this work, we focused on generative models and selected mT5 (Xue et al., 2021) and XGLM Lin et al. (2021) due to their broad language coverage, highly multilingual training data, and pre-training with up-sampling of low-resource languages to ensure more balanced language representation. However, many other models are available. For instance, BLOOM (Le Scao et al., 2023) is a decoder-only model trained on 46 languages, although low factual recall accuracy has been reported (Qi et al., 2023). Llama-2 (Touvron et al., 2023) has been used in some multilingual analyses (Wendler et al., 2024; Dumas et al., 2024), but it is primarily an English-centric model given that 90% of its pre-training data is in English. More recent models that could have been considered include mGPT (Shliazhko et al., 2022) and EuroLLM (Martins et al., 2024), both decoder-only models trained on 61 and 34 languages, respectively. On the other hand, the Aya-101 model (Üstün et al., 2024) is an instruction fine-tuned version of mT5; however, we focus on pre-trained multilingual LMs to compare our analysis to the English monolingual pre-trained versions previously studied.
Appendix B Causal Tracing
@@
Figure 9: XGLM causal analysis for each language (rest of the languages in Figure 10). Average probability (of the originally predicted token) recovered when corrupting the subject in the input and restoring the hidden representation at a given layer (State), or restoring the MLP / Self Attn. in a window of 10 layers.
@@
Figure 10: XGLM causal analysis for each language (Rest of the languages in Figure 9). Average probability (of the originally predicted token) recovered when corrupting the subject in the input and restoring the hidden representation at a given layer (State), or restoring the MLP / Self Attn. in a window of 10 layers.
@@
Figure 11: Causal analysis of mT5 for each language (Rest of the languages in Figure 12). Average probability (of the originally predicted token) recovered when corrupting the subject in the input and restoring the hidden representation at a given layer (State), or restoring the MLP / Self Attn. / Cross Attn. in a window of 6 layers.
@@
Figure 12: Causal analysis of mT5 for each language (Rest of the languages in Figure 11). Average probability (of the originally predicted token) recovered when corrupting the subject in the input and restoring the hidden representation at a given layer (State), or restoring the MLP / Self Attn. / Cross Attn. in a window of 6 layers.
Appendix C Attention Knockout
@@
Figure 14: Attention knockout for each language in mT5. We present the results from knocking out the information from the last token in the decoder (last) and the masked object in the encoder (enc_sentinel). However, from the encoder sentinel there is no much flow of information so these are not plot in the main body of the paper.
Appendix D Extraction Event
@@
Figure 17: Number of extraction events split by precedence (or not) of an extraction event in the self-attn (cross-attn). Left XGLM, right mT5.
@@
Appendix E Patching
XGLM mT5
$=$ Relation, $\neq$ Subject $\neq$ Relation, = Subject $=$ Relation, $\neq$ Subject $\neq$ Relation, = Subject
Source - Target All $o_{s}^{l_{s}}\neq o_{s}^{l_{t}}$ $o_{t}^{l_{s}}\neq o_{t}^{l_{t}}$ All $o_{s}^{l_{s}}\neq o_{s}^{l_{t}}$ $o_{t}^{l_{s}}\neq o_{t}^{l_{t}}$ All $o_{s}^{l_{s}}\neq o_{s}^{l_{t}}$ $o_{t}^{l_{s}}\neq o_{t}^{l_{t}}$ All $o_{s}^{l_{s}}\neq o_{s}^{l_{t}}$ $o_{t}^{l_{s}}\neq o_{t}^{l_{t}}$
en-es 849 155 140 325 75 82 842 830 779 263 250 249
en-vi 862 369 377 337 177 180 854 844 823 209 207 204
en-tr 974 260 253 68 36 15 815 802 794 209 206 205
en-ru 823 565 588 218 187 177 796 795 787 191 191 190
en-uk 914 734 715 35 32 32 853 850 852 136 136 133
en-ko 995 993 991 6 6 6 744 744 743 165 165 165
en-he 962 903 458 0 0 0 778 778 778 166 166 166
en-fa 537 528 513 0 0 0 714 714 712 114 114 114
en-ar 829 818 816 212 209 211 775 775 774 159 159 159
Table 2: Total number of pairs of examples considered in the patching experiments.
XGLM mT5
$=$ Relation, $\neq$ Subject $\neq$ Relation, $=$ Subject $=$ Relation, $\neq$ Subject $\neq$ Relation, $=$ Subject
Patch - Context $\mathcal{L}_{c}(o_{r,s_{p}})$ $\mathcal{L}_{p}(o_{r,s_{c}})$ $\mathcal{L}_{c}(o_{r_{p},s})$ $\mathcal{L}_{p}(o_{r_{c},s})$ $\mathcal{L}_{c}(o_{r,s_{p}})$ $\mathcal{L}_{p}(o_{r,s_{c}})$ $\mathcal{L}_{c}(o_{r_{p},s})$ $\mathcal{L}_{p}(o_{r_{c},s})$
en-es 1.3% (2) 38.6% (54) 49.3% (37) 7.3% (6) 0.7% (6) 21.6% (168) 0.8% (2) 10.4% (26)
en-vi 0.5% (2) 67.1% (253) 11.3% (20) 13.3% (24) 0.2% (2) 20.4% (168) 1.0% (2) 13.7% (28)
en-tr 1.2% (3) 52.6% (133) 47.2% (17) 33.3% (5) 0.2% (2) 26.3% (209) 2.4% (5) 11.2% (23)
en-ru 2.3% (13) 53.1% (312) 63.6% (119) 6.2% (11) 1.4% (11) 36.1% (284) 7.9% (15) 11.6% (22)
en-uk 1.1% (8) 26.2% (187) 15.6% (5) 18.8% (6) 0.4% (3) 31.8% (271) 4.4% (6) 9.0% (12)
en-ko 0.4% (4) 45.0% (446) 83.3% (5) 33.3% (2) 0.5% (4) 32.3% (240) 4.2% (7) 13.3% (22)
en-he 0.0% (0) 57.9% (265) - - 0.3% (2) 38.4% (299) 5.4% (9) 16.3% (27)
en-fa 0.0% (0) 42.3% (217) - - 0.8% (6) 25.3% (180) 3.5% (4) 16.7% (19)
en-ar 8.8% (72) 39.7% (324) 40.7% (85) 4.7% (10) 0.9% (7) 32.0% (248) 2.5% (4) 9.4% (15)

### v2

@@ 6 Conclusion
In this paper, we analyzed factual knowledge recall mechanisms in 10 languages using multilingual transformer-based LMs, comparing them to prior research on English recall in monolingual autoregressive LLMs. We discovered that some mechanisms, such as the flow of subject representations in the later layers and the extraction phase, are present in multilingual and monolingual models. However, we also identified notable differences, including the localization of knowledge in not only one but two MLP sites; and the joint role of MLPs and attention modules during the extraction phase in multilingual models. A key contribution of our work is the first-ever investigation of language encoding during recall, achieved through patching representations. In decoder-only models, the relation flows to the last token first, followed by the language. In contrast, in mT5, both relation and language flow to the last token at similar layers. This suggests that while relation and subject representations are multilingual and enable cross-lingual object extraction, the extraction phase itself is language-specific, as language encoding precedes extraction. These findings provide new evidence to understand the factual knowledge recall in transformer LMs, and to how decoder-only LMs resolve tasks in stages. Contributing with new directions for the study of cross-lingual transfer and knowledge localization.
@@ 7 Limitations
In this paper, we examined three model architectures, leaving out the effects of model sizes, instruction fine-tuning, or models like Llama that can behave multilingually but have less coverage and less multilingual pre-training data. Additionally, our analyses were conducted on 500-1000 examples per language, which we believe provides a sufficient sample size for generalization; however, the results are inherently limited by the relations present in the mParaRel dataset, which may not capture all factual nuances. Additionally, although we analyzed 10 diverse languages, many more languages exist, and further research is needed to confirm the generalizability of our findings across a broader linguistic spectrum. Lastly, we described the main mechanisms found in XGLM, EuroLLM and mT5, however other weaker mechanisms could be at play, which could describe, for example, the low extraction rates found for some languages (Figure 27) or the few examples where the object seems to be encoded from early layers before the extraction takes place (Figure 47).
@@ Appendix A Related Work
Factual Knowledge Recall
Petroni et al. (2019) analyzed factual knowledge in pre-trained language models, by creating the LAMA dataset, which contains natural language templates and WikiData triplets. Subsequently, Elazar et al. (2021) developed the ParaRel dataset, which contains manually paraphrases of the LAMA templates, to measure the consistency of factual knowledge.  [^15] One may object to calling the ability to solve cloze tests as knowledge. LLMs are sometimes inconsistent and cannot always justify the propositions they generate. We side with Fierro et al. (2024) in thinking knowledge may nevertheless be the appropriate concept, since LLMs are mostly consistent and can sometimes provide justification, e.g., in virtue of world models or training data attribution. This led to further research in multilingual settings, Kassner et al. (2021); Jiang et al. (2020) translated LAMA, creating mLAMA and X-FACTR, and evaluated the factual knowledge acquired by multilingual pre-trained models, while Wang et al. (2024a) evaluated knowledge from the perspective of knowledge edits and their impact between Chinese and English (Zhu et al., 2021). Building on this, Yin et al. (2022) curated a dataset that focuses on the cultural aspect of factual knowledge. While Fierro and Søgaard (2022) studied multilingual consistency of knowledge using mParaRel, a machine-translated version of the English templates in ParaRel, where subjects and objects are translated using WikiData identifiers. In a related effort, Qi et al. (2023) measured the cross-lingual consistency of multilingual LLMs by evaluating the consistency across languages, rather than just across paraphrases in the same language. Qi et al. (2023) used the mLAMA dataset. In this paper, we used mParaRel since it is an extended version of mLAMA, as it is a translation of the paraphrases in mParaRel.
Interpretability on Factual Knowledge Recall
@@
Activation Patching
Ghandeharioun et al. (2024) introduced the Patchscope framework to study what an intermediate representation encodes. In their experiments, they focused on “decoding” the contents of a latent representation by patching it into the forward pass of another input. More closely related to our work, Dumas et al. (2024) patched the last token representation to investigate if models encode language-agnostic concepts in a word translation task. Their findings, consistent with ours and those of Foroutan et al. (2022), support the existence of language-agnostic representations. Similarly, Wang et al. (2024b) patched the last token representation to study how relations are encoded in factual English queries, finding that the last token encodes the query’s relation at a specific computational stage—aligning with our results. In this paper, we are able to draw more general conclusions given that we study a cross-lingual factual recall setup. While Dumas et al. (2024) concluded that LLMs first resolve the output language, we find that the model first resolves the query’s relation and then the language. Thus, we interpret the last token representation as a Function Vector (FV) (Todd et al., 2024), encoding the function to solve the given task: for Dumas et al. (2024) the task was translation so the function encodes the language to which translate. Finally, we are also able to show that for decoder-only models the FV composes the language onto the relation FV, which extends the findings from Wang et al. (2024b).
@@
In this work, we focused on generative models and selected mT5 (Xue et al., 2021), XGLM Lin et al. (2021), and EuroLLM (Martins et al., 2024) due to their broad language coverage, highly multilingual training data, and pre-training with up-sampling of low-resource languages to ensure more balanced language representation. However, many other models are available. For instance, BLOOM (Le Scao et al., 2023) is a decoder-only model trained on 46 languages, although low factual recall accuracy has been reported (Qi et al., 2023). Llama-2 (Touvron et al., 2023) has been used in some multilingual analyses (Wendler et al., 2024; Dumas et al., 2024), but it is primarily an English-centric model given that 90% of its pre-training data is in English. More recent models that could have been considered include mGPT (Shliazhko et al., 2022), a decoder-only model trained on 61. On the other hand, the Aya-101 model (Üstün et al., 2024) is an instruction fine-tuned version of mT5; however, we focus on pre-trained multilingual LMs to compare our analysis to the English monolingual pre-trained versions previously studied.
Language Script Family SOV/SVO
English Latin Germanic SVO
Spanish Latin Romance SVO
Vietnamese Latin Austroasiatic SVO
Turkish Latin Turkic SOV
Russian Cyrillic Slavic SVO*
Ukrainian Cyrillic Slavic SVO*
Japanese Kanji Proto-Japonic SOV
Korean Korean Koreanic SOV
Hebrew Hebrew Arabic VSO
Farsi (Persian) Perso-Arabic Indo-Iranian SOV
Arabic Arabic Arabic VSO
Table 2: Languages, their scripts, families, and sentence structures (SVO: subject-verb-object, SOV: subject-object-verb, VSO: verb-subject-object, SVO*: SVO dominant but SOV is also possible).
Appendix B Experimental Setup
We use the mParaRel triplets and templates to query the factual knowledge of the language models. As mentioned earlier, we perform 3 modifications to the dataset. First, we fetch WikiData for aliases of the target object to be able to match different possible surface forms. Second, we filter out examples where the target object is contained in the query, e.g. Microsoft Outlook is developed by. Finally, to better compare across languages we control the variety of the subject-object pairs, by only using a crosslingual version of mParaRel. Specifically, for each relation we filter out triplets that are not present in all the languages (in mParaRel a subject and object may have not been translated if they were not found in WikiData). Thus, in the crosslingual mParaRel version, each subject-object pair in a relation is present in each of the languages. For XGLM we additionally restrict the templates to have the object at the end of the sentence, therefore for some languages (tr, ko, ja, and fa) both the autorregressive condition on the templates and the crosslingual restriction leads to too few total examples (< 1000), so for these we do not impose the crosslingual restriction. For all the other languages, and for all the languages in mT5 there are enough examples so we restrict them to be crosslingual. In Table 3 we present the total number of examples we consider per language and model, and the correctly predicted fraction. [^16] In decoder-only models, the total number of examples is higher in some languages due to crosslingual filtering, which is applied per relation. If a language has no examples for a given relation, it doesn’t restrict examples in other languages. Since decoder-only models use only autoregressive templates, some lower-resource languages may have zero examples for certain relations. However, when using all templates (as in mT5), these lower-resource languages can restrict the triplets available for other languages within the same relation.
XGLM and EuroLLM mT5
Language Correct Total Percentage Correct Total Percentage
en 1812 4147 43.7% 1543 3853 40.0%
es 1380 4167 33.1% 1192 3926 30.4%
vi 1646 4068 40.5% 993 3748 26.5%
tr 418 2278 18.3% 1058 4033 26.2%
ru 830 4133 20.1% 683 3826 17.9%
uk 213 4144 5.1% 456 3830 11.9%
ko 116 1444 8.0% 630 3783 16.7%
ja 6 1790 0.3% 358 4124 8.7%
he 13 4403 0.3% 565 4064 13.9%
fa 7 4388 0.2% 406 3814 10.6%
ar 811 4482 18.1% 488 4110 11.9%
Table 3: Total number of examples in mParaRel.
B.1 Prompt Details
For mT5, we feed the mParaRel input into the encoder with a sentinel token in the object placeholder. In the decoder, we provide the beginning-of-sequence token followed by the sentinel token. We only check the tokens generated next to the first sentinel token, as the pre-training task of the model is to generate the text for each sentinel in the input, the decoder usually continues to generate answers for other sentinel tokens. Any tokens generated preceding the object, are added to the decoder input since adding these to the encoder does not ensure that the next token predicted will be the object.
B.2 Computational Resources
The experiments were run in A100 GPUs. The causal analysis took from 30 minutes to 12 hours depending on the language. The rest of the experiments took 1-2 hours per language.
Appendix C Causal Tracing
C.1 XGLM
@@
Figure 9: XGLM causal analysis for each language (continues in Figure 10). Average logit score (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 4 layers, or the Self Attn. in a window of 4 layers.
@@
Figure 10: XGLM causal analysis for each language (Rest of the languages in Figure 9). Average logit score (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 4 layers, or the Self Attn. in a window of 4 layers.
@@
Figure 11: XGLM causal analysis for each language (continues in Figure 12). Average probability (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 4 layers, or the Self Attn. in a window of 4 layers.
@@
Figure 12: XGLM causal analysis for each language (Rest of the languages in Figure 11). Average probability (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 4 layers, or the Self Attn. in a window of 4 layers.
C.2 EuroLLM
Figure 13: EuroLLM causal analysis for each language (continues in Figure 14). Average logit score (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 5 layers, or the Self Attn. in a window of 5 layers.
Figure 14: EuroLLM causal analysis for each language (Rest of the languages in Figure 13). Average logit score (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 5 layers, or the Self Attn. in a window of 5 layers.
Figure 15: EuroLLM causal analysis for each language (continues in Figure 16). Average probability (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 5 layers, or the Self Attn. in a window of 5 layers.
Figure 16: mT5 causal analysis for each language (Rest of the languages in Figure 15). Average probability (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 5 layers, or the Self Attn. in a window of 5 layers.
C.3 mT5
Figure 17: mT5 causal analysis for each language (continues in Figure 18). Average logit score (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 3 layers, or the Self Attn. in a window of 3 layers.
Figure 18: mT5 causal analysis for each language (Rest of the languages in Figure 17). Average logit score (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 3 layers, or the Self Attn. in a window of 3 layers.
Figure 19: mT5 causal analysis for each language (continues in Figure 20). Average probability (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 3 layers, or the Self Attn. in a window of 3 layers.
Figure 20: mT5 causal analysis for each language (Rest of the languages in Figure 19). Average probability (for the originally predicted token) recovered after corrupting the subject in the input and restoring: the hidden representation at a given layer (State), the MLP in a window of 3 layers, or the Self Attn. in a window of 3 layers.
Appendix D Attention Knockout
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 21: Attention knockout between the last token and a given set of tokens. Each layer represents the effect of the knockout on a window of $w$ layers. Models from top to bottom: XGLM ($w=6$), EuroLLM ($w=7$), mT5 ($w=4$).
@@ Appendix D Attention Knockout
Figure 23: EuroLLM attention knockout for each language.
Figure 24: Attention knockout for each language in mT5. The knockout is performed between the last token in the decoder (“last”) to a given set of tokens $\{t\}$, and between the masked token in the encoder (“enc_sentinel”) and $\{t\}$. From the encoder sentinel there is no much flow of information so these were not included in the main body.
Appendix E Extraction Event
@@
Figure 26: EuroLLM extraction rates for each language.
@@
Figure 28: Number of extraction events split by precedence (or not) of an extraction event in the self-attn or cross-attn. Models from top to bottom: XGLM, EuroLLM, mT5.
@@
Figure 30: Number of extraction events split by precedence (or not) of an extraction event in the self-attn in EuroLLM for each language.
@@
Appendix F Patching
XGLM EuroLLM mT5
Patch - Context All $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ All $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ All $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$
en-es 849 834 759 880 862 755 842 831 780
en-vi 862 857 836 930 910 843 854 844 825
en-tr 974 961 961 966 945 914 815 803 794
en-ru 823 821 817 846 846 827 796 795 787
en-uk 914 914 911 915 915 915 853 850 852
en-ko 995 995 991 991 991 991 744 744 743
en-he 962 962 961 810 810 775 778 778 778
en-fa 537 537 532 771 771 734 714 714 712
en-ar 829 829 828 862 859 858 775 775 774
en-ja 134 134 134 895 895 872 774 774 773
Table 4: Total number of patch-context examples considered in the patching experiments with $\{\neq\mathcal{L},=r,\neq s\}$. The $\mathcal{L}_{c}(o_{p})$ column is the total number of examples where the detection of $\mathcal{L}_{c}(o_{p})$ would be unambiguous, that is, $\mathcal{L}_{c}(o_{p})\neq\mathcal{L}_{p}(o_{p})$, conversely for the $\mathcal{L}_{p}(o_{c})$ column.
XGLM EuroLLM mT5
Patch - Context $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$
en-es 0.7% (6) 26.0% (197) 5.0% (43) 34.2% (258) 0.5% (4) 21.5% (168)
en-vi 0.5% (4) 34.4% (288) 0.5% (5) 16.1% (136) 0.1% (1) 20.2% (167)
en-tr 0.2% (2) 16.4% (158) 0.2% (2) 25.4% (232) 0.4% (3) 26.2% (208)
en-ru 2.2% (18) 51.4% (420) 4.6% (39) 63.5% (525) 1.1% (9) 36.1% (284)
en-uk 0.5% (5) 21.3% (194) 6.9% (63) 62.5% (572) 0.2% (2) 31.8% (271)
en-ko 0.3% (3) 47.1% (467) 0.1% (1) 75.7% (750) 0.3% (2) 32.3% (240)
en-he 0.0% (0) 27.3% (262) 0.0% (0) 65.9% (511) 0.4% (3) 38.4% (299)
en-fa 0.0% (0) 41.7% (222) 0.3% (2) 64.0% (470) 1.0% (7) 25.1% (179)
en-ar 8.7% (72) 39.5% (327) 4.5% (39) 65.7% (564) 0.6% (5) 31.9% (247)
en-ja 0.0% (0) 3.0% (4) 1.9% (17) 11.5% (100) 0.0% (0) 26.0% (201)